# How Far Are We From AGI?

### v1

@@ 2.3 AI Memory
Current State of AI Memory We introduce this part in terms of memory management and utilization:
  • AI Memory management. Memory can be categorized in two ways: duration and representation. In terms of duration, memory is classified into short-term and long-term memory. Regarding representation, memory is divided into textual memory and parametric memory.
@@
      Long-term memory: Long-term memory can be broadly classified into two main types: experiences and knowledge. Experiences encompass a range of elements such as past observations, thoughts, actions, and more. These can be documented and analyzed through various forms such as history event flows (Park et al., 2023), agent trajectories (Yao et al., 2020), thought trajectories (Yao et al., 2023), feedback summaries (Shinn et al., 2023), past input-output pairs (Rubin et al., 2021), etc. This rich collection of experiences serves a critical function in decision-making processes. By retrieving relevant experiences, agents can gain additional information necessary for reasoned judgment, understand feedback from past actions, and achieve a level of generalization in their understanding and reasoning.
@@
      Knowledge represents an agent’s understanding of the world and itself, which enhances its reasoning and decision-making capabilities. Knowledge can originate from two sources. First, AI agents can collect and assimilate knowledge from experiences, integrating newly acquired information or skills into their existing knowledge. For example, Voyager (Wang et al., 2023h) maintains a memory that includes executable codes for preliminary actions to accomplish tasks. Second, AI agents or models can utilize external knowledge bases. For instance, ReAct (Yao et al., 2022) employs Wikipedia APIs to acquire external knowledge when agents lack information during their activities. ChatGPT Browse with Bing enables ChatGPT to access internet knowledge for answering questions, significantly enhancing its ability to provide accurate responses (OpenAI, 2023a). Retrieval-augmented methods (Lewis et al., 2020; Guu et al., 2020; Shuster et al., 2021; Borgeaud et al., 2022) leverage a knowledge base of unstructured text. The “reading to learn" methods (Branavan et al., 2012; Hanjie et al., 2021) utilize domain knowledge from text manuals to influence the policies in reinforcement learning.
    3. 3.
@@
      Textual memory vs. parametric memory: Textual memory is the prevalent method for representing memory content today. It can include both unstructured formats like raw natural language and structured forms such as tuples, databases, and more. Alternatively, memory can be represented in a parametric form. Techniques like supervised fine-tuning (Hu et al., 2021) or knowledge editing (De Cao et al., 2021; Mitchell et al., 2021) can integrate domain-specific knowledge into model parameters. For textual memory, each inference involves incorporating memory into the context prompt, leading to higher costs and extended processing times during the reading and inference processes. Conversely, parametric memory often incurs greater costs during the writing phase, as fine-tuning models is more challenging than simple text storage. Regarding interpretability, textual memory is generally more transparent than parametric memory, as the natural language provides the most direct means for human understanding (Zhang et al., 2024a).
@@
  • AI Memory utilization. There are two common technologies to utilize memories: memory retrieval and long-context LLMs.
@@
      Memory retrieval: Memory retrieval involves reading information from long-term memory to short-term memory for immediate use. Various agents retrieve experiences to guide their behavior. Reflexion (Shinn et al., 2023) reflects on task feedback signals and maintains them as textual summaries. These summaries are directly incorporated into the context of subsequent episodes, aiding in performance enhancement. Voyager (Wang et al., 2023h) maintains a continuously expanding skill library of executable code for storing complex behaviors, allowing for embedding-based retrieval in similar future situations. Generative agents (Park et al., 2023) document their experiences in natural language and retrieve memories using a mix of relevance (embedding-based), recency (rule-based), and importance (reasoning-based) criteria.
      Retrieval-augmented methods enhance language generation by retrieving knowledge. The Dense Passage Retriever (DPR) (Karpukhin et al., 2020) creates dense representations of documents and retrieves the most relevant documents based on their prior probability using Maximum Inner Product Search (MIPS). The Retrieval-Augmented Language Model pre-training (REALM) (Guu et al., 2020) integrates unsupervised pre-training of a knowledge retriever with masked language modeling, enabling direct retrieval of documents to supplement language predictions. Retrieval-Augmented Generation (RAG) models (Lewis et al., 2020; Shuster et al., 2021; Borgeaud et al., 2022) employ a non-parametric memory, such as a dense vector index of Wikipedia, accessed via a pre-trained neural retriever (e.g., DPR). These documents are processed by a seq2seq model, which conditions its output generation on both the input and the retrieved documents. Both the retriever and seq2seq modules, initialized from pre-trained models, are jointly fine-tuned, allowing both retrieval and generation to adapt to downstream tasks.
@@
      Long-context LLMs: The expansion of the context window in long-context LLMs or LVLMs opens up new avenues for models to access their long-term memory. This has been made possible by more powerful GPUs with enhanced memory capabilities and breakthroughs in memory-efficient attention mechanisms (Dao et al., 2022; Tay et al., 2022), allowing the context window for pre-trained LLMs to increase from 1024 tokens in GPT-2 (Radford et al., 2019), to 8192 in GPT-4 (Achiam et al., 2023), and now exceeding 16K tokens. With these expanded context windows, AI systems can more effectively store and recall knowledge and experiences within their context, enabling faster and more comprehensive context-based reasoning.
@@
  • Future AGI will enhance memory utilization through the integration of retrieval and advanced reasoning, enabling more human-like intelligence and adaptability. Beyond simple memory retrieval, future AGI systems could enhance memory utilization by integrating retrieval procedures with advanced reasoning capabilities. This involves recalling information and synthesizing and applying it in strategic and context-appropriate ways. Beyond fixed implementations, the retrieval procedures could also be learned or updated to adapt to changing circumstances. The ability to access and apply relevant information from their memory in real-time would be a significant step towards more human-like intelligence, enabling these systems to respond to new situations with a high level of understanding and adaptability.
  • Future AGI will autonomously update their knowledge, enabling continuous learning and adaptation while ensuring safety. Unlike existing retrieval-augmented models that primarily rely on pre-existing, human-generated content, future AGI systems could autonomously generate, evaluate, and incorporate new content into their memory banks. These updates should encompass knowledge essential for performance enhancement and experiences the systems can draw upon. This concept is closely linked to self-evolve, which we will discuss later. It would allow AGIs to learn from their own experiences and insights, continually enriching and updating their knowledge base. In a constantly changing world, this capability will also enable the systems to adapt quickly given new information and unlearn outdated knowledge. A crucial aspect is to guarantee the safety of the memory updates, ensuring that no harmful information is written that could lead to contamination.
@@ 2.4 AI Metacognition
I am no bird; and no net ensnares me: I am a free human being with an independent will.
@@
Metacognition represents a suite of advanced cognitive and emotional capacities critical for organizational effectiveness, encompassing the nuanced interpretation of complex situations, deep self-awareness, emotional resilience under pressure, and the intrinsic motivation to lead and innovate. These abilities are vital for the articulation of tacit knowledge and fostering an environment conducive to organizational learning, thereby serving as key drivers for both personal and institutional growth (Choudrie and Selamat, 2006).
Current State of AI Metacognition The discourse on metacognition extends into the realm of AGI, where such capabilities are deemed equally critical. For AGI systems, metacognition, such as self-awareness, consciousness, the capacity for self-evolution, and theory of mind, are posited to be foundational for bridging the gap towards achieving genuine general intelligence. These internal abilities are pivotal in enabling AGI systems to autonomously learn, efficiently complete tasks, and align more closely with human intentions.
  • Self-Awareness in AGI. Developing self-awareness in AI, particularly within the realm of robotics, hinges on intricate concepts such as self-reflection, meta-cognition, and different self-perception techniques, including self-distancing and self-immersion. These concepts are integral to constructing social robots equipped with cognitive architectures that support self-description, the utilization of personal pronouns, and the ability to respond to self-focusing cues, which are fundamental for facilitating effective human interactions and environmental navigation. As AI systems evolve, the philosophical and practical considerations of equipping them with human-like traits of conscientiousness are gaining traction, heralding a burgeoning field of research (Huang et al., 2023c). For a thorough understanding of this area, interdisciplinary approaches that intertwine psychology, artificial intelligence, and ethics are instrumental.
@@
  • AGI holding certain persona. Recent advancements reveal that LLMs can exhibit consistent personality traits, such as those categorized by the Big Five or MBTI frameworks, with models like ChatGPT often exhibiting traits aligned with the ENFJ type (Huang et al., 2023c). These models also tend to display certain cognitive thinking styles, with evidence suggesting an inclination towards holistic thinking in ChatGPT’s responses (Jin et al., 2023b). Research efforts are increasingly directed towards intentionally imbuing LLMs with specific personalities, enabling them to demonstrate a variety of behaviors that are both diverse and verifiable (Caron and Srivastava, 2022; Jiang et al., 2022b).
    These traits are crucial for various reasons. First, self-awareness could enhance AGI’s adaptive problem-solving abilities by allowing it to accurately assess its strengths and limitations, thus facilitating adjusting its strategies in real-time when faced with new challenges. Second, the capacity for ethical and moral decision-making is increasingly imperative as AGI becomes more entwined with societal functions, necessitating a self-awareness component to enable navigation through complex moral dilemmas and ensure alignment with human values. Furthermore, the potential for AGI to autonomously evolve and adapt without human guidance promises greater efficiency and capability in the long term, possibly leading to an exponential increase in their abilities—a key characteristic of true AGI. Lastly, incorporating autonomous consciousness in AGI could yield more natural and effective human-AI interactions, enhancing collaboration and developing more intuitive user interfaces, particularly in scenarios requiring deep integration into human teams or societal structures.
  • AGI metacognition ability in self-evolving. While the aforementioned research defines AGI in terms of easily measurable capabilities (Butlin et al., 2023; Morris et al., 2024), it may overlook the potential importance of cognitive abilities such as self-evolution or self-awareness. Studies predominantly showcase this through the agent’s iterative adaptation via task execution (Le, 2019; Wang et al., 2023e), code execution (Gao et al., 2020), or feedback from physical simulations (Qian et al., 2024; Xu et al., 2023). Other strategies for self-evolution include prompt adaptation and optimization (Wang et al., 2023h; Aksitov et al., 2023), continuous improvement through error identification and self-reflection, and memory retrieval as a mechanism for short- or long-term learning. These approaches mainly emphasize the iterative refinement of tasks within a loop-structured framework based on LLMs. In contrast, recent advancements propose methodologies that address inter-task agent self-evolution, highlighting the significance of leveraging past experiences to effectively evolve AI systems (Qian et al., 2024; Xu et al., 2023).
@@ 3 AGI Interface: Connecting the World with AGI
In the pursuit of developing AGI, a crucial aspect to address is its capability to interact with the external world. This interaction is facilitated through various interfaces that enable AGI systems to perceive, understand, and act within their environment, be it digital, physical, or intellectual. We summarize these thire future directions in Figure 4.
@@ 3.1 AI Interfaces to Digital World
The concept of AGI interface into the digital world extends the scope by allowing agents to interact with digital environments, such as the Internet, databases, code, and APIs, and exhibit intelligent behaviors similar to human-like behavior (Qin et al., 2023a). Interface to Digital World provides an accessible, practical, and scalable testbed for studying advanced cognitive skills in agents, promoting their development through language-based inputs and the ease of developing digital tools.
@@ 3.2 AI Interfaces to Physical World
The integration of AI into physical entities is a crucial aspect of the pursuit of AGI. AGI in the physical world emphasizes learning through direct interaction with the environment and making an impact on reality, such as creating or modified substances. In this section, we will explore the latest advancements in embodied AI in the physical world, including robotic control, navigation, and manipulation.
Current State of AI Interfaces to the Physical World The current state mainly lies in the interaction with robotic functionalities, understanding the potential for more intuitive human-robot interfaces and emphasizing the importance of real-world datasets in advancing AI’s practical applications.
  • Robotic control and action. Recent advancements in robotic control and action demonstrate the potential for robots to interpret and execute complex, high-level instructions through natural language, providing a more intuitive interface between humans and robots. SayCan (Ahn et al., 2022) combines the semantic understanding capabilities of Google’s PaLM with robotic affordances, enabling robots to understand abstract tasks and execute them in real-world environments. PaLM-E (Driess et al., 2023) injects embodied observations into the language embedding space of a pre-trained language model for sequential robotic manipulation planning, while VIMA (Jiang et al., 2022a) processes multi-modal prompts and outputs motor actions autoregressively to control a robot arm. RT2 (Brohan et al., 2023) leverages large-scale, pre-trained vision-language models for robotic control.
@@
AGI-level Interfaces to the Physical World Looking ahead, the journey toward AGI with embodied intelligence encompasses several key areas for future research and development. A critical area is enhancing contextual and environmental understanding in AI systems. The ability to interpret and adapt to dynamic, real-world environments in a manner akin to human cognition.

### v2

@@ 2.3 AI Memory
Current State of AI Memory We examine the current state of AI memory, focusing on three key aspects: memory management, which determines what and when to store; memory representation, which defines how information is structured; and memory utilization, which addresses how to apply and use the memory efficiently and effectively.
  • Memory management. Memory is categorized by duration into short-term and long-term memory.
@@
      Long-term memory: Long-term memory can be broadly classified into two main types: experiences and knowledge. Experiences encompass a range of elements such as past observations, thoughts, actions, and more. This rich collection of experiences serves a critical function in decision-making processes. By retrieving relevant experiences, agents can gain additional information necessary for reasoned judgment, understand feedback from past actions, and achieve a level of generalization in their understanding and reasoning. For example, Reflexion (Shinn et al., 2023) reflects on task feedback signals and maintains them as textual summaries. These summaries are directly incorporated into the context of subsequent episodes, aiding in performance enhancement. Generative agents (Park et al., 2023) document their experiences in natural language and retrieve memories using a mix of relevance (embedding-based), recency (rule-based), and importance (reasoning-based) criteria.
@@
      Knowledge represents an agent’s understanding of the world and itself, which enhances its reasoning and decision-making capabilities. Knowledge can originate from two sources. First, AI agents can collect and assimilate knowledge from experiences, integrating new information or skills into their existing knowledge. For example, Voyager (Wang et al., 2023h) maintains a continuously expanding skill library of executable codes for preliminary actions to accomplish tasks. Second, AI agents or models can utilize external knowledge bases. For instance, ReAct (Yao et al., 2022) employs Wikipedia APIs to acquire external knowledge when agents lack information during their activities. ChatGPT Browse with Bing enables ChatGPT to access internet knowledge for answering questions, significantly enhancing its ability to provide accurate responses (OpenAI, 2023a). Retrieval-augmented methods (Lewis et al., 2020; Guu et al., 2020; Shuster et al., 2021; Borgeaud et al., 2022) leverage a knowledge base of unstructured text. The “reading to learn” methods (Branavan et al., 2012; Hanjie et al., 2021) utilize domain knowledge from text manuals to influence the policies in reinforcement learning.
@@
  • Memory representation. Regarding representation, memory is divided into textual memory and parametric memory. Textual memory is the prevalent method for representing memory content today. It can include both unstructured formats like raw natural language and structured forms such as tuples, databases, and more. Alternatively, memory can be represented in a parametric form. Techniques like supervised fine-tuning (Hu et al., 2021), knowledge editing (De Cao et al., 2021; Mitchell et al., 2021) and model merging (Du et al., 2024; Yang et al., 2024; Lu et al., 2024; Goddard et al., 2024) can integrate domain-specific knowledge into model parameters. For textual memory, each inference involves incorporating memory into the context prompt, leading to higher costs and extended processing times during the reading and inference processes. Conversely, parametric memory often incurs greater costs during the writing phase, as fine-tuning models is more challenging than simple text storage. Regarding interpretability, textual memory is generally more transparent than parametric memory, as the natural language provides the most direct means for human understanding (Zhang et al., 2024a).
@@
  • Memory utilization. There are two common technologies to utilize memories: memory retrieval and long-context LLMs.
@@
      Memory retrieval: Memory retrieval involves reading information from long-term memory to short-term memory for immediate use. This can be accomplished through rule-based retrieval or retrieval-augmented methods. Rule-based retrieval can search memory using keywords, timesteps, or specific patterns. In retrieval-augmented approaches, the Dense Passage Retriever (DPR) (Karpukhin et al., 2020) creates dense representations of documents and retrieves the most relevant documents based on their prior probability using Maximum Inner Product Search (MIPS). The Retrieval-Augmented Language Model pre-training (REALM) (Guu et al., 2020) integrates unsupervised pre-training of a knowledge retriever with masked language modeling, enabling direct retrieval of documents to supplement language predictions. Retrieval-Augmented Generation (RAG) models (Lewis et al., 2020; Shuster et al., 2021; Borgeaud et al., 2022) employ a non-parametric memory, such as a dense vector index of Wikipedia, accessed via a pre-trained neural retriever (e.g., DPR). These documents are processed by a seq2seq model, which conditions its output generation on both the input and the retrieved documents. Both the retriever and seq2seq modules, initialized from pre-trained models, are jointly fine-tuned, allowing both retrieval and generation to adapt to downstream tasks.
@@
      Long-context LLMs: The expansion of the context window in long-context LLMs opens up new avenues for models to access their long-term memory. Works like Ring Attention (Liu et al., 2023j) and LongRoPE (Ding et al., 2024b) greatly reduce the time and cost of long context inference by improving the operation mechanism and storage method of attention. More powerful GPUs with enhanced memory capabilities and further breakthroughs in memory-efficient attention mechanisms (Dao et al., 2022; Tay et al., 2022), allowing the context window for pre-trained LLMs to increase from 1024 tokens in GPT-2 (Radford et al., 2019), to 8192 in GPT-4 (Achiam et al., 2023), and now exceeding 16K tokens. With these expanded context windows, AI systems can more effectively store and recall knowledge and experiences within their context, enabling faster and more comprehensive context-based reasoning.
@@
  • Future AGI will enhance memory utilization through the integration of retrieval and advanced reasoning, enabling more human-like intelligence and adaptability. Beyond simple memory retrieval, future AGI systems could refine memory utilization by intricately combining retrieval processes with advanced reasoning that strategically synthesizes and applies information in context-appropriate ways. Beyond fixed implementations, the retrieval procedures could be learned or updated to adapt to changing circumstances. The ability to access and apply relevant information from their memory in real-time would be a significant step towards more human-like intelligence, enabling these systems to respond to new situations with a high level of understanding and adaptability.
  • Future AGI will autonomously update their knowledge, enabling continuous learning and adaptation while ensuring safety. Unlike existing retrieval-augmented models that primarily rely on pre-existing, human-generated content, future AGI systems could autonomously generate, evaluate, and incorporate new content into their memory banks. These updates should encompass knowledge essential for performance enhancement and experiences the systems can draw upon. This concept is closely linked to self-evolve, which we will discuss later. It would allow AGIs to learn from their own experiences and insights, continually enriching and updating their knowledge base. In a constantly changing world, this capability will also enable the systems to adapt quickly given new information and unlearn outdated knowledge. A crucial aspect is to guarantee the safety of the memory updates, ensuring that no harmful information is written that could lead to contamination. Designing safety constraints for autonomous AGIs involves creating robust validation protocols that assess the truthfulness, relevance, and impact of new information before integration. We can implement expert systems to periodically review updates, use anomaly detection to flag outliers and potentially harmful data, and employ additional methods to enforce these safety constraints.
@@ 2.4 AI Metacognition
I am no bird, and no net ensnares me: I am a free human being with an independent will.
@@
Metacognition (Choudrie and Selamat, 2006) of humans involves key cognitive and emotional skills such as understanding complex situations, self-awareness, and motivation to innovate. These abilities help share implicit knowledge and drive personal growth.
Current State of AI Metacognition The discourse on metacognition extends into the realm of AGI, where such capabilities are deemed equally critical. For AGI systems, metacognition, such as self-awareness (Chella et al., 2020; Subagdja et al., 2021), consciousness (Dehaene et al., 2021), the capacity for self-evolution (Floreano et al., 2004; Tao et al., 2024), and theory of mind (Cuzzolin et al., 2020), are posited to be foundational for bridging the gap towards achieving AGI. These internal abilities can enable AI systems to autonomously learn, efficiently complete tasks, and align more closely with human intentions.
  • Self-Awareness in AGI. Developing self-awareness in AI, particularly within the realm of robotics (Scassellati, 2002), hinges on intricate concepts such as self-reflection (Shinn et al., 2024), meta-cognition (Langdon et al., 2022), and self-distancing (Kross and Ayduk, 2017). These concepts are integral to constructing social robots equipped with cognitive architectures that support self-description, the utilization of personal pronouns, and the ability to respond to self-focusing cues, which are fundamental for facilitating effective human interactions and environmental navigation. As AI systems evolve, the philosophical and practical considerations of equipping them with human-like traits of conscientiousness are gaining traction, heralding a burgeoning field of research (Huang et al., 2023c). For a thorough understanding of this area, interdisciplinary approaches that intertwine psychology, artificial intelligence, and ethics are instrumental.
@@
  • AGI holding certain persona. Recent advancements reveal that LLMs can exhibit consistent personality traits, such as those categorized by the Big Five or MBTI frameworks, with models like ChatGPT often exhibiting traits aligned with the ENFJ type (Huang et al., 2023c). These models also tend to display certain cognitive thinking styles, with evidence suggesting an inclination towards holistic thinking in ChatGPT’s responses (Jin et al., 2023c). Research efforts are increasingly directed towards intentionally imbuing LLMs with specific personalities, enabling them to demonstrate a variety of behaviors that are both diverse and verifiable (Caron and Srivastava, 2022; Jiang et al., 2022b).
  • AGI metacognition ability in self-evolving. While the aforementioned research defines AGI in terms of easily measurable capabilities such as reasoning (Butlin et al., 2023; Morris et al., 2024), it may overlook the potential importance of meta-cognitive abilities such as self-evolution or self-awareness. Studies predominantly showcase this through the agent’s iterative adaptation via task execution (Le, 2019; Wang et al., 2023e), code execution (Gao et al., 2020), or feedback from physical simulations (Qian et al., 2024; Xu et al., 2023a). Other strategies for self-evolution include prompt adaptation and optimization (Wang et al., 2023h; Aksitov et al., 2023), continuous improvement through error identification and self-reflection, and memory retrieval as a mechanism for short- or long-term learning. These approaches mainly emphasize the iterative refinement of tasks within a loop-structured framework based on LLMs. In contrast, recent advancements propose methodologies that address inter-task agent self-evolution, highlighting the significance of leveraging past experiences to effectively evolve AI systems (Qian et al., 2024; Xu et al., 2023a).
These traits are crucial for various reasons. First, self-awareness could enhance AGI’s adaptive problem-solving abilities by allowing it to accurately assess its strengths and limitations, thus facilitating adjusting its strategies in real time when faced with new challenges. Second, the capacity for ethical and moral decision-making is increasingly imperative as AGI becomes more entwined with societal functions, necessitating a self-awareness component to enable navigation through complex moral dilemmas and ensure alignment with human values. Furthermore, the potential for AGI to autonomously evolve and adapt without human guidance promises greater efficiency and capability in the long term, possibly leading to an exponential increase in their abilities—a key characteristic of true AGI. Lastly, incorporating autonomous consciousness in AGI could yield more natural and effective human-AI interactions, enhancing collaboration and developing more intuitive user interfaces, particularly in scenarios requiring deep integration into human teams or societal structures.
@@ 3 AGI Interface: Connecting the World with AGI
In the pursuit of developing AGI, a crucial aspect to address is its capability to interact with the external world. This interaction is facilitated through various interfaces that enable AGI systems to perceive, understand, and act within their environment, be it digital, physical, or intellectual. We summarize these three future directions in Figure 4.
@@ 3.1 AI Interfaces to Digital World
The concept of AGI interface into the digital world extends the scope by allowing agents to interact with digital environments, such as the Internet, databases, code, and APIs, and exhibit intelligent behaviors similar to human-like behavior (Qin et al., 2023a). This interface serves as a crucial bridge for grounding AGI in complex, real-world scenarios, providing an indispensable platform for simulating and interacting with the multifaceted nature of human knowledge and experience. By facilitating AGI’s engagement with real-world information structures and problem-solving contexts, this digital world interface accelerates the development of more versatile and robust artificial general intelligence capable of operating effectively across various domains.
@@ 3.2 AI Interfaces to Physical World
The integration of AI into physical entities is a crucial aspect of the pursuit of AGI. AGI in the physical world emphasizes learning through direct interaction with the environment and making an impact on reality, such as creating or modifying substances. In this section, we will explore the latest advancements in embodied AI in the physical world, including robotic control, navigation, and manipulation.
Current State of AI Interfaces to the Physical World The current state mainly lies in the interaction with robotic functionalities, understanding the potential for more intuitive human-robot interfaces, and emphasizing the importance of real-world datasets in advancing AI’s practical applications.
  • Robotic control and action. Recent advancements in robotic control and action including PaLM-E (Driess et al., 2023), RT-2 (Zitkovich et al., 2023), and Mobile Aloha (Fu et al., 2024b) demonstrate the potential for robots to interpret and execute complex, high-level instructions through natural language, providing a more intuitive interface between humans and robots. SayCan (Ahn et al., 2022) combines the semantic understanding capabilities of PaLM (Chowdhery et al., 2022a) with robotic affordances, enabling robots to understand abstract tasks and execute them in real-world environments. PaLM-E injects embodied observations into the language embedding space of a pre-trained language model for sequential robotic manipulation planning, while VIMA (Jiang et al., 2022a) processes multi-modal prompts and outputs motor actions autoregressively to control a robot arm. RT2 leverages large-scale, pre-trained vision-language models for robotic control.
@@
AGI-level Interfaces to the Physical World Looking ahead, the journey toward AGI with embodied intelligence encompasses several key areas for future research and development. A critical area is enhancing contextual and environmental understanding in AI systems. AGI needs to develop the capacity to interpret and adapt to dynamic, real-world environments with a level of sophistication at least comparable to human cognition. This advanced environmental analysis and perception capability is essential for AGI to effectively navigate and interact with complex, ever-changing physical surroundings, mirroring the adaptability and intuition that humans exhibit in diverse situations.